OpenStudy (anonymous):

Find the GCF of 12, 18, and 84 A. 6 B. 12 C. 216 D. 252

OpenStudy (anonymous):

6 12/6=2 18/6=3 84/6=14

OpenStudy (anonymous):

I also go 6 :)